OpenStudy (linda):

express in radical form : (sin pi over 4) (tan pi over 6) - (tan pi over 3) (cos pi over 2) plz help

OpenStudy (anonymous):

dang what happened to my answer?

OpenStudy (anonymous):

ok second term is 0 since \[cos(\frac{\pi}{2})=0\]so we can ignore it

OpenStudy (anonymous):

first term is \[sin(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

and \[tan(\frac{\pi}{6})=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

multiplying gives \[\frac{\sqrt{2}}{2\times \sqrt{3}}\]

OpenStudy (anonymous):

to write in simplest radical form multiply top and bottom by \[\sqrt{3}\] to get \[\frac{\sqrt{6}}{6}\]

OpenStudy (linda):

ok thats it?